OpenStudy (anonymous):

x^2+y=120 find y. what is the max area? (the problem is this: a rectangle and a fence to cover 3 sides, find area of the longest side.)

OpenStudy (anonymous):

Is the Perimeter = 120?

OpenStudy (anonymous):

WELL THE FENCING NEED FOR 3 SIDES IS 120

OpenStudy (anonymous):

o.k. then 2x + y = 120 because two of the sides are the same, 2x, and the other side doesn't have to be, y. so using this equation you get y = 120 - 2x Area = Length * Width = y * x = (120 - 2x)*x That is A = 120x - 2x^2 or prettier: \[A = 120x -2x^2\] This is an upside down parabola so the maximum will be at the vertex. When you find the x value of the vertex, plug it into the equation \[A = 120x -2x^2\] to get the maximum area Does that help?